Bhadbabyjay03:

Consider this linear function: y=1/2x+1 Plot all ordered pairs for the values in the domain. D: {-8, -4, 0, 2, 6}

Mercury:

the domain is the set of x-values in the function so, you'd need to find the corresponding y-value for each x-value, then plot the (x,y) values let's start with the first value, x = -8 plug this into the function y=1/2x+1 y= (1/2)(-8) + 1 = -4 + 1 = -3 so you'd plot (-8,-3) as your first point repeat with the rest of the domain values